Wellplace Insurance Company processes insurance policy applications in batches of 100. One day, they had 10 batches to process and, after inspection, it was found that four batches had nonconforming policies. One batch had 3 nonconformances, another had 5, another had 2, and another had 1 nonconformance. What were (a) the proportion nonconforming for each batch, (b) the nonconformances per unit (NPU), in total for the 10 batches, and (c) the total throughput yield (TY) for the 10 batches?
Answer #1

Total number of policies = 100 policies/ batch x 10 batches = 1000 policies

Proportion of non conforming for each batch :

Batch number

Batch size

Number of non conformances

Proportion of non conformances

1 to 6

100

0

0

7

100

3

0.03

8

100

5

0.05

9

100

2

0.02

10

100

1

0.01

Total number of policies in 10 batches = 1000

Total number of non conformances in 10 batches = 3 _+ 5 + 2 + 1 = 11

Thus, non conformance per unit for 10 batches = 11/1000 = 0.011

NON CONF0RMANCE PER UNIT IN TOTAL FOR TEN BATCHES = 0.011

Correct and conforming policies = 1000 – 11 = 989

Therefore, total throughput yield for 10 batches = 989/1000 x 100 = 98.9%

TOTAL THROUGHPUT YIELD FOR 10 BATCHES = 98.9%